Home‎ > ‎Crew 164‎ > ‎

0227-DailySummaryReport

DAILY SUMMARY REPORT
SOL: 6
Person filling out Report: Mohammad IRANMANESH, XO
Summary Title:  Mars is not trying to kill us anymore.
Mission Status: All systems GO. 

Sol Activity Summary: Started working on a mid rotation video. Atmosphere opacity experiment is working but weather condition not optimal for test. Rover range test. EVA. Work on projects. Change of water filter with a 3D printed spare. We started tracking internet consumption hour per hour to avoid any mistake.

Look Ahead Plan:  Finish mid-rotation video. Rest tomorrow afternoon before starting next week experiments.

Anomalies in work: None
Weather: Hot but cloudy. 
EVA: ATV worked well. Scouted North and South of the Hab mainly staying on the main road. No big issue.

Reports to be filed: 
- Commander report
- Engineering report
- Journalist report 
- Astronomy report
- EVA #6 report
- EVA #7 request

Support Requested: /